# Towards a unified hadron-quark equation of state for neutron stars within the relativistic mean-field model

### v1

@@ I Introduction
A common strategy for constructing a hybrid EoS is to “paste” independent models for the hadronic and quark phases. While sophisticated descriptions exist for each sector individually, this two-phase construction introduces an artificial separation and requires additional assumptions to connect the two regimes. Most notably, the hadron–quark transition is typically imposed through a Maxwell or Gibbs construction, rather than arising directly from the underlying theory. An alternative and more physically robust strategy is to develop a unified framework in which both hadrons and quarks are described within a single effective Lagrangian. In such models, like the chiral mean-field (CMF) approach [4, 5], inspired by the Polyakov-loop extended Nambu–Jona-Lasinio (PNJL) model [6], a scalar field $\Phi$ is introduced to mediate deconfinement, dynamically suppressing quark degrees of freedom at low densities and temperatures.
Motivated by this approach, we introduce a similar $\Phi$-dependent mechanism into the density-dependent relativistic mean-field (DDRMF) parametrization SW4L [7, 8, 9], leading to a unified hadron–quark model named EVA–01, the first Extended VAriable-coupling relativistic mean-field model that unifies hadronic and quark phases within a single Lagrangian. This construction allows for a consistent description of both sectors, preserving the nuclear phenomenology of the hadronic interaction while extending it to include quark degrees of freedom within the same field-theoretical structure.
@@ III.2 Astrophysics
An important clarification regarding the inclusion of the crust for the outer layers must be made. EVA-01 is not capable of describing the behavior of matter in the low-density region of the crust. For this reason, independent models are used, and the interface between the crust and hadronic is set to occur at $n=0.5\,n_{0}$. Furthermore, we adopt different crust EoSs to reflect the physical expectation that the crust cools monotonically, even as the core temporarily heats up [92, 89, 93]. Specifically, we use the hot crust EoS from Dehman et al. [94] with $T=5$ MeV for stage $1$ and $T=2$ MeV for stage $2$. For the final cold configuration (stage $3$), we employ the traditional BPS-BBP crust [95, 96]; although more modern models exist, the traditional BPS–BBP EoS is still widely used, as it remains consistent with current knowledge in the low-density regime (for more details regarding the use of the BPS-BBP crust and a comparison with other crust models see, for example, Refs. [97, 98] and references therein)It is important to clarify that the crust EoSs employed for the three stages are almost identical in the outer crust region, as Dehman et al. [94] used a $T=0$ approximation in this sector; in the inner crust, where finite temperature effects become relevant, the EoS crust used for the hot stage $1$ is the stiffest, that for the cold stage $3$ is the softest, and the hot stage $2$ EoS crust lies in between for most of the inner crust density range.

### v2

@@ I Introduction
A common strategy for constructing a hybrid EoS is to “paste” independent models for the hadronic and quark phases. While sophisticated descriptions exist for each sector individually, this two-phase construction introduces an artificial separation and requires additional assumptions to connect the two regimes. Most notably, the hadron–quark transition is typically imposed through a Maxwell or Gibbs construction, rather than arising directly from the underlying theory. An alternative and more physically robust strategy is to develop a unified framework in which both hadrons and quarks are described within a single effective Lagrangian. Recently, unified descriptions based on quarkyonic matter and parity-doublet dynamics have been proposed, in which quark degrees of freedom coexist with partial confined matter at intermediate densities [4]. In other models, like the chiral mean-field (CMF) approach [5, 6], inspired by the Polyakov-loop extended Nambu–Jona-Lasinio (PNJL) model [7], a scalar field $\Phi$ is introduced to mediate deconfinement, dynamically suppressing quark degrees of freedom at low densities and temperatures.
Motivated by this last approach, we introduce a similar $\Phi$-dependent mechanism into the density-dependent relativistic mean-field (DDRMF) parametrization SW4L [8, 9, 10], leading to a unified hadron–quark model named EVA–01, the first Extended VAriable-coupling relativistic mean-field model that unifies hadronic and quark phases within a single Lagrangian. This construction allows for a consistent description of both sectors, preserving the nuclear phenomenology of the hadronic interaction while extending it to include quark degrees of freedom within the same field-theoretical structure.
@@ III.2 Astrophysics
An important clarification regarding the inclusion of the crust for the outer layers must be made. EVA-01 is not capable of describing the behavior of matter in the low-density region of the crust. For this reason, independent models are used, and the interface between the crust and hadronic is set to occur at $n=0.5\,n_{0}$. Furthermore, we adopt different crust EoSs to reflect the physical expectation that the crust cools monotonically, even as the core temporarily heats up [93, 90, 94]. Specifically, we use the hot crust EoS from Dehman et al. [95] with $T=5$ MeV for stage $1$ and $T=2$ MeV for stage $2$. For the final cold configuration (stage $3$), we employ the traditional BPS-BBP crust [96, 97]; although more modern models exist, the traditional BPS–BBP EoS is still widely used, as it remains consistent with current knowledge in the low-density regime (for more details regarding the use of the BPS-BBP crust and a comparison with other crust models see, for example, Refs. [98, 99] and references therein). It is important to clarify that the crust EoSs employed for the three stages are almost identical in the outer crust region, as Dehman et al. [95] used a $T=0$ approximation in this sector; in the inner crust, where finite temperature effects become relevant, the EoS crust used for the hot stage $1$ is the stiffest, that for the cold stage $3$ is the softest, and the hot stage $2$ EoS crust lies in between for most of the inner crust density range.